A chocolate chip cookie manufacturing company recorded the number of chocolate chips in a sample of 70 cookies. The mean is 22.2422.24 and the standard deviation is 2.442.44.Construct a 98​% confidence interval estimate of the standard deviation of the numbers of chocolate chips in all such cookies.

How do I find the degree of freedom, which is 69, using the chi critical value table? 69 is not listed on there, just 60 and 70. Not sure how to use my TI-84 calculator to figure it out.
Answer #1

At alpha = 0.02,

ZCritical = 2.33

Mean = 22.24

Std Dev = 2.44

n = 70

Std Error = 2.44/701/2 =0.292

98% CI = Mean +/- ZCritical * Std Error = 22.24 +/- 2.33*0.292 = {21.56,22.92}

Since you are looking for 69 and table contains 70, you can use 70 as it is nearest to your requirement.